1. On the planet Gizmo, the inhabitants travel by high speed trains that run on air tracks much like the air track you used in lab. A train car with a mass of 8700 kg is traveling at 11.0 m/s when it strikes a second car moving in the same direction at 2.2 m/s. The two stick together and move off with a speed of 4.00 m/s. What is the mass of the second car? Explain your reasoning.

2. How much kinetic energy was lost in the collision?
Answer #1

1) Ans: 6.6 m/s

According to coservarcons of momentum

Initial Momentum = Final Momentum

m1u1 + m2u2. = (m1+m2)v

8700×11 + 8700×2.2 = (8700+8700)v

114840 = 17400×v

⇒ v = 6.6 m/s

2) Ans: 16 J

∆E = | (K.E)intial - (K.E)final |

. = | ( m1u12/2 + m2u22 ) - (m1+m2)v2/2 |

∆E = | 8700×112/2 + 8700×2.22/2 - 17400×6.62/2 |

= | 547404 - 547420 |

⇒ ∆E = | - 16 | J